Maria has twice as many quarters as dimes if she has 7.00 dollars how many of each type of coin does she have ?
Answer by math_helper(2461)

Maria has twice as many quarters as dimes if she has 7.00 dollars how many of each type of coin does she have ?
—————————————————————————————————
Number of dimes: d
Number of quarters: 2d
25(2d) + 10(d) = 700 (in cents)
60d = 700
d = 700/60 = 11.66667
There seems to be a mistake/omission in the problem statement. She can not have a fractional number of dimes and quarters.
( If she had $7.20 then the problem would have an integer solution (d=12, 2d=24) )